a charge q is placed at the center of a circle of a cube of side l.what is the electric flux passing through each face of the cube

Grade:12

1 Answers

Arun
5 years ago

According to Gauss’ Theorem, the total electric flux through all the six faces of the cube is,

φ = Charge enclosed / ε 

φ = q /ε 

Each face of the cube comprises 1/6 of the total surface of the cube. Therefore, the flux passing through each face is,

φ = 1/6 ? q/ε  =  q /6 ε 0